### v1

@@ 4.2 Transfer Performance of AutoGraph
We evaluate the transferability of AutoGraph by pre-training it on a large dataset of synthetic graphs generated using NetworkX (Hagberg et al., 2008) and fine-tuning it on the unattributed graph datasets. Dataset and experimental details are provided in Appendix E. As shown in Table 8, the pre-trained model consistently outperforms the baseline on small synthetic datasets in terms of the VUN score, achieving near-perfect validity. On larger graph datasets, the pre-trained model also surpasses the baseline across MMD metrics, demonstrating its ability to generalize to more complex structures. However, on small synthetic datasets, the pre-trained model shows a slight decline in MMD metrics compared to the baseline. These findings highlight the potential of building foundation models for graph generation and underscore the need for more comprehensive benchmark datasets, similar to those established in other domains.
Table 8: Transfer performance on downstream tasks using AutoGraph pre-trained on the NetworkX dataset. Red and green colors indicate relative decreases and increases respectively, compared to AutoGraph without pre-training.
Dataset Deg. Clus. Orbit Spec. Ratio VUN (improv.)
@@ 4.3 Substructure Conditioned Generation
We explore the ability of AutoGraph to perform substructure-conditioned generation without requiring fine-tuning. Given a subgraph $S$ (which could represent a functional motif of interest in drug discovery), we flatten the subgraph into a SENT sequence and condition the generation process on this sequence. This approach guarantees that the generated graph will contain $S$ as an induced subgraph (Thm. 2.11). As a proof-of-concept, we follow the methodology of Vignac et al. (2023); Maziarz et al. (2022) and generate molecular graphs starting from a specific motif, called 1,4-Dihydroquinoline [^1] https://pubchem.ncbi.nlm.nih.gov/compound/1_4-Dihydroquinoline, using the model pre-trained on the GuacaMol dataset. Our results in Table 9 demonstrate that this approach maintains similar validity, uniqueness, and novelty to unconditional generation (Table 7). To further showcase the flexibility of this method, we test more extreme cases by replicating the same motif multiple times before performing the conditional generation. While validity decreases significantly when using an unrealistically large number of copies (e.g, 5), the model still generates some visually plausible molecules (Appendix F.3), showing superior flexibility over Vignac et al. (2023). These results highlight the potential of AutoGraph for important applications in drug discovery, particularly in motif scaffolding.
Table 9: Motif scaffolding using the motif 1,4-Dihydroquinoline.
# Copies of the motif Valid Unique Novelty
@@ 4.4 Ablation Experiments
In this study, we aim to understand the effectiveness of key components in AutoGraph.
@@
AutoGraph provides a novel framework for evaluating the capability of current LLM architectures in graph generation and, more broadly, in structural reasoning tasks. In Table 10, we compare several state-of-the-art architectures on the Planar dataset, including GPT-2 (Radford et al., 2019), Mamba (Gu & Dao, 2023), and LLaMA (Touvron et al., 2023a). While all models achieve comparable MMD ratios, transformer-based architectures, particularly LLaMA, demonstrate significantly better performance in terms of VUN scores compared to state-space models. These findings highlight the potential of AutoGraph to serve as a valuable benchmark for assessing sequence/language models’ capabilities in graph generation tasks.
Table 10: Comparison of sequence model architectures on Planar
@@
A key advantage of AutoGraph over diffusion-based approaches is the flexibility to apply top-k sampling (Fan et al., 2018) during inference, which can improve generation quality. As shown in Figure 2, a smaller $k$ improves the VUN score on the Planar dataset, whereas it is not beneficial on the SBM dataset. In contrast, increasing $k$ generally improves MMD ratios across both datasets. These observations suggest that top-k sampling can be optimized based on dataset characteristics. In our experiments, we select the best $k$ that maximizes the VUN score for small synthetic datasets and minimizes the validation MMD ratios for other datasets. Importantly, this flexibility allows practitioners to select $k$ based on the specific performance criteria they aim to prioritize.
Figure 2: The effect of top-k sampling on the Planar and SBM datasets.
@@
As discussed in Section 2, SENT is preferred over SET for graph generation, as incorporating neighborhood information is essential to ensure structural coherence. To empirically validate this, we compare the performance of SENT and SET on the Planar dataset and present the training curves in Figure 3. Consistent with our theoretical analysis, SET fails to produce high-validity graphs, resulting in a VUN score close to zero, whereas SENT successfully generates valid planar graphs.
Figure 3: SET vs SENT on the Planar dataset
@@ 5 Conclusion
We proposed AutoGraph, a scalable and efficient autoregressive model for attributed graph generation. AutoGraph demonstrates the ability to handle large graphs while preserving high generation quality. It enables substructure-conditioned generation without requiring additional fine-tuning and shows promising transfer capabilities. More importantly, it establishes a critical connection between graph modeling and language modeling, representing a significant step toward leveraging language modeling paradigms to address challenges in graph generation or more general graph learning tasks.
@@ Acknowledgements
Impact Statement
@@
Appendix
This appendix provides both theoretical and experimental materials. It is organized as follows: Section A provides additional background on sequence models. Section B provides information about model availability and limitations. Section C provides additional details and remarks on our method. Section D provides proofs for the theorems presented in the main manuscript. Section E provides experimental details. Section F provides additional quantitative and qualitative results.
Appendix A Background on Sequence Model Architectures
Our sequence model architectures are fully based on established natural language models. In particular, we consider three prominent models, including GPT-2 (Radford et al., 2019), LLaMA (Touvron et al., 2023a; b), and Mamba (Gu & Dao, 2023) to demonstrate the effectiveness of our approach. Notably, our methodology is not restricted to these specific models; it can be applied to any sequence or language model.
GPT-2.
GPT-2 represents one of the earliest large language models based on the transformer architecture (Vaswani et al., 2017). The model employs pre-normalization with LayerNorm, the GeLU activation function, and absolute positional embeddings to encode token positions in sequences. These design choices laid the foundation for many subsequent models.
LLaMA.
LLaMA (Touvron et al., 2023a; b) builds upon the transformer framework with several key enhancements. It incorporates pre-normalization through RMSNorm (Zhang & Sennrich, 2019) and employs the SwiGLU activation function (Shazeer, 2020). Additionally, LLaMA replaces absolute positional embeddings with rotary positional embeddings (Su et al., 2024), enabling better generalization to longer sequences.
Mamba.
Mamba (Gu & Dao, 2023) is a state-space model (SSM) that maps input sequences to outputs using continuous-time dynamics. It introduces a selection mechanism that dynamically controls how input data flows into hidden states, making the model parameters adaptive to time and data. This innovation enables Mamba to achieve superior performance compared to other SSMs across various tasks.
Appendix B Availability and Limitations
B.1 Code and Model Availability
Our code and all pre-trained models will be released upon publication.
B.2 Limitations
@@ C.1 Background on Graph Theory
We provide additional background on graph theory necessary for the definitions and theories of SETs and SENTs. The background is largely based on Diestel (2005).
We first give the formal definition of graph isomorphism:
@@
Graph isomorphism is an equivalence relation on graphs and as such it partitions the class of all graphs into equivalence classes. A set of graphs isomorphic to each other is called an isomorphism class of graphs. It is worth noting that our SENT isomorphism also partitions the class of all SENTs into equivalence classes in a similar fashion.
We also provide the formal definition of induced subgraph:
@@ C.2 Remarks on Tokenized SENTs
We now detail how to modify the original graph $G$: we introduce three virtual nodes, labeled /, $\bm{<}$, and $\bm{>}$ respectively. These virtual nodes are connected to all other virtual nodes and original nodes in the graph. This modification ensures that for any non-special token in the tokenized sequence, its subsequent token can either be one of its neighbors or one of the virtual nodes (/, $\bm{<}$, or $\bm{>}$). Consequently, each token has a direct connection to the node corresponding to the current token, and the language model amounts to learning the state transition functions for these random walks. Since these random walks are non-Markovian, this perspective further justifies our choice of using autoregressive models instead of one-step generative models. Furthermore, as random walks are random sequences on graphs, sampling random walks amounts to sampling from those random sequences.
@@ C.3 Remarks on Model Inference
The model inference is straightforward following the same process as LLMs such as LLaMA (Touvron et al., 2023a; b). An alternative way is to enforce the semantic correctness of the generated sequences of tokens by adjusting the logits at a certain token to obey the semantic rule of the tokenization. For instance, the token ‘$\bm{>}$’ can only occur after a token ‘$\bm{<}$’ or no special tokens can appear right after /. We manually implemented these transition constraints and incorporated them into the inference. We compared this strategy with the constraint-free counterpart. Surprisingly, our experiments demonstrate that the constraint-free variant could always generate semantically correct tokenized SENTs and perform similarly to the one with the transition constraints. Therefore, we did not use any transition constraints during the inference in our experiments.
@@ Appendix D Proofs
[…]
@@ E.1 Datasets
Both of these datasets are from Martinkus et al. (2022). The Planar dataset consists of 200 planar graphs with 64 nodes each, generated via Delaunay triangulation on points uniformly sampled in the unit square. The SBM dataset contains 200 graphs comprising 2 to 5 communities, with each community having between 20 and 40 nodes. An edge is placed between two nodes with probability 0.3 if they belong to the same community, and 0.05 otherwise. We follow the same splits as Martinkus et al. (2022).
@@
The Proteins dataset includes graph representations (contact maps) of proteins from Dobson & Doig (2003). In these graphs, each node represents an amino acid, and an edge connects two nodes if their corresponding amino acids are within 6 angstroms of each other. We use the same data splits as Liao et al. (2019). The Point Clouds dataset, also from Liao et al. (2019), consists of 41 point clouds of household objects (Neumann et al., 2013). As many of these graphs are disconnected, we retain only the largest connected component of each, following Bergmeister et al. (2024), and again employ the splits used by Liao et al. (2019).
@@
The QM9 dataset, from Wu et al. (2018), comprises small molecules with up to nine heavy atoms (carbon, oxygen, nitrogen, and fluorine). In this work, we adopt the more challenging setting proposed by Vignac et al. (2023), where hydrogen atoms are modeled explicitly, and we follow the same data splits as in that reference.
@@
The MOSES and GuacaMol datasets are obtained from the respective benchmark tools of Polykovskiy et al. (2020) and Brown et al. (2019). Both consist of drug-like molecules, with those in GuacaMol typically being larger on average. For each dataset, we convert generated molecular graphs to SMILES using the code from Jo et al. (2022), which permits partial charges. We employ the standard data splits provided by the corresponding benchmarks.
@@
We generate the graphs using the generators from the NetworkX library [^2] https://networkx.org/documentation/stable/reference/generators.html (Hagberg et al., 2008), categories including “Classic”, “Lattice”, “Small”, “Random Graphs”, “Geometric”, “Trees”, “Community”, “Social Networks”. We ensure that this dataset does not contain any graphs in the downstream datasets. The summary of the code for generating these graphs is provided in Table 12. Notably, the largest graph has up to 5999 nodes.
@@ E.2 Evaluation Metrics
We follow Martinkus et al. (2022) and Vignac et al. (2023) in comparing our model’s performance with other graph generative approaches. Specifically, we measure the maximum mean discrepancy (MMD) between the generated and test graphs for degree distribution, clustering coefficient, orbit counts, and spectrum. As a reference, we also compute these metrics on the training set and report the mean ratio across all properties as a global indicator of statistical discrepancy between the generated samples and test samples. Note that for the Point Clouds dataset, which is defined by a $k$-nearest-neighbor structure, the degree MMD is always zero and is therefore excluded from the mean ratio. While we utilize these metrics to maintain consistency with previous research, we acknowledge their limitations, particularly regarding arbitrary kernel hyperparameter selection, as highlighted by O’Bray et al. (2022); Thompson et al. (2022).
We additionally track uniqueness and novelty: uniqueness is the fraction of generated graphs that are not isomorphic to each other, and novelty is the fraction of generated graphs that are not isomorphic to any training graph.
Below, we describe additional metrics specific to each dataset.
@@
Following Martinkus et al. (2022), we report a validity score for synthetic datasets. For Planar graphs, it verifies whether the generated graphs remain planar; for SBM graphs, it measures how likely they are to be generated under the original SBM parameters.
@@
For QM9, we report the validity, uniqueness, and novelty defined for general molecules, as described in the following paragraph. We also report atom stability and molecule stability as defined by Hoogeboom et al. (2022) and Vignac et al. (2023).
@@
Since MOSES (Polykovskiy et al., 2020) and GuacaMol (Brown et al., 2019) are benchmarking platforms, each comes with its own suite of metrics, which we use to evaluate our model. These include:
@@ E.4 Hyperparameters
Unlike prior studies that adjust model sizes across datasets, we maintain a consistent model architecture and size throughout all experiments, specifically using the small GPT configuration (768 hidden dimensions, 12 layers, 12 attention heads). Training hyperparameters are aligned with established practices from popular LLMs such as GPT-3 (Brown et al., 2020) and LLaMA (Touvron et al., 2023a). We fix the context length to 2048 and use a batch size of 128 if possible, otherwise 64 for larger graphs. In particular, we employ the AdamW optimizer with a gradient clipping threshold of 1.0, a weight decay of 0.1, and a learning rate schedule with a linear warmup followed by cosine decay, peaking at 6e-4. The AdamW hyperparameters are set to $\beta=(0.9,0.95)$. Due to the small dataset sizes of previous benchmarks, we tune the only training hyperparameter dropout in $\{0,0.5\}$, and find the model achieves better validation loss with the value of 0.5 on the small synthetic datasets.
Inference hyperparameters, including $k$ (top-k sampling) and $\tau$ (temperature), are reported in Table 13 and analyzed in detail in Section 4.4 in the main manuscript.

### v2

@@ 4.2 Transfer Performance of AutoGraph
We evaluate the transferability of AutoGraph by pre-training it on a large dataset of synthetic graphs generated using NetworkX [23] and fine-tuning it on the unattributed graph datasets. Dataset and experimental details are provided in Appendix E. As shown in Table 5, the pre-trained model consistently outperforms the baseline on small synthetic datasets in terms of the VUN score, achieving near-perfect validity. On larger graph datasets, the pre-trained model also surpasses the baseline across MMD metrics, demonstrating its ability to generalize to more complex structures. However, on small synthetic datasets, the pre-trained model shows a slight decline in MMD metrics compared to the baseline. These findings highlight the potential of building foundation models for graph generation and underscore the need for more comprehensive benchmarks beyond synthetic datasets.We also test the transferability of AutoGraph on molecular graphs, by pre-training it on the PubChem-10M dataset [12] and fine-tuning on GuacaMol. The pre-trained model substantially outperforms the baseline, as shown in Table 4.
Table 5: Transfer performance on downstream tasks using AutoGraph pre-trained on the NetworkX dataset. Red and green colors indicate relative decreases and increases, respectively, compared to AutoGraph without pre-training.
Dataset Deg. Clus. Orbit Spec. Ratio VUN (improvement)
@@ 4.3 Substructure Conditioned Generation
Table 6: Motif scaffolding
# Motif copies Valid Unique Novelty
@@
We explore the ability of AutoGraph to perform substructure-conditioned generation without requiring fine-tuning. Given a subgraph $S$ (which could represent a functional motif of interest in drug discovery), we flatten the subgraph into a SENT sequence and condition the generation process on this sequence. This approach guarantees that the generated graph will contain $S$ as an induced subgraph (Thm. 2.14). As a proof-of-concept, we follow the methodology of Vignac et al. [61], Maziarz et al. [42] and generate molecular graphs starting from a specific motif, called 1,4-Dihydroquinoline [^1] https://pubchem.ncbi.nlm.nih.gov/compound/1_4-Dihydroquinoline, using the model pre-trained on the GuacaMol dataset. Our results in Table 6 demonstrate that this approach maintains similar validity, uniqueness, and novelty to unconditional generation (Table 4). To further showcase the flexibility of this method, we test more extreme cases by replicating the same motif multiple times before performing the conditional generation. While validity decreases significantly when using an unrealistically large number of copies (e.g, 5), the model still generates some visually plausible molecules (Appendix F.4), showing superior flexibility over Vignac et al. [61]. These results highlight the potential of AutoGraph for important applications in drug discovery, particularly in motif scaffolding. Additional experiments on multiple but different motifs are provided in App. F.4.
@@ 4.4 Ablation Experiments
In this study, we aim to understand the effectiveness of the key components in AutoGraph.
@@
AutoGraph provides a novel framework for evaluating the capability of current LLM architectures in graph generation and, more broadly, in structural reasoning tasks. In Table 7, we compare several state-of-the-art architectures on the Planar dataset, including GPT-2 [49], Mamba [22], and LLaMA [58]. While all models achieve comparable MMD ratios, transformer-based architectures, particularly LLaMA, demonstrate significantly better performance in terms of VUN scores compared to state-space models. These findings highlight the potential of AutoGraph to serve as a valuable benchmark for assessing sequence/language models’ capabilities in graph generation tasks.
Table 7: Comparison of sequence model architectures on the Planar dataset.
@@
A key advantage of AutoGraph over diffusion-based approaches is the flexibility to apply top-k sampling [19] during inference, which can improve generation quality. As shown on the left of Figure 2, a smaller $k$ improves the VUN score on the Planar dataset, whereas it is not beneficial on the SBM dataset. In contrast, increasing $k$ generally improves MMD ratios on both datasets. These results suggest that top-k sampling can be optimized based on dataset characteristics. In our experiments, we select the best $k$ that maximizes the VUN score for small synthetic datasets and minimizes the validation MMD ratios for other datasets. This flexibility allows practitioners to select $k$ based on specific performance criteria they aim to prioritize.
@@
As discussed in Section 2, SENT is preferred over SET for graph generation, as incorporating neighborhood information is essential to ensure structural coherence. To empirically validate this, we compare the performance of SENT and SET on the Planar dataset and present the training curves on the right of Figure 2. Consistent with our theoretical analysis, SET fails to produce high-validity graphs, resulting in a VUN score close to zero, whereas SENT successfully generates valid planar graphs.
Figure 2: Ablation experiments. Left: the effect of top-k sampling on the Planar and SBM datasets. Right: the validation loss and VUN scores when using SET and SENT on the Planar dataset.
@@ 5 Conclusion
We proposed AutoGraph, a scalable and efficient autoregressive model for attributed graph generation that handles large graphs while maintaining high quality. Our approach enables substructure-conditioned generation without additional fine-tuning and demonstrates promising transfer capabilities. Crucially, AutoGraph establishes the first fundamental connection between graph and language modeling–where graphs are losslessly represented as token sequences, and prefixes in these sequences serve as meaningful patterns in both paradigms–representing a significant step toward applying language modeling techniques to graph generation and broader graph learning challenges.
Limitations.
@@ Acknowledgements
Appendix
This appendix provides both theoretical and experimental materials. It is organized as follows: Section A provides additional background on sequence models. Section B provides a discussion about the broader impact of this work. Section C provides additional details and remarks on our method. Section D provides proofs for the theorems presented in the main manuscript. Section E provides experimental details. Section F provides additional quantitative and qualitative results.
Appendix A Background on Sequence Model Architectures
Our sequence model architectures are fully based on established natural language models. In particular, we consider three prominent models, including GPT-2 [49], LLaMA [58, 59], and Mamba [22] to demonstrate the effectiveness of our approach. Notably, our methodology is not restricted to these specific models; it can be applied to any sequence or language model.
GPT-2.
GPT-2 represents one of the earliest large language models based on the transformer architecture [60]. The model employs pre-normalization with LayerNorm, the GeLU activation function, and absolute positional embeddings to encode token positions in sequences. These design choices laid the foundation for many subsequent models.
LLaMA.
LLaMA [58, 59] builds upon the transformer framework with several key enhancements. It incorporates pre-normalization through RMSNorm [67] and employs the SwiGLU activation function [51]. Additionally, LLaMA replaces absolute positional embeddings with rotary positional embeddings [54], enabling better generalization to longer sequences.
Mamba.
Mamba [22] is a state-space model (SSM) that maps input sequences to outputs using continuous-time dynamics. It introduces a selection mechanism that dynamically controls how input data flows into hidden states, making the model parameters adaptive to time and data. This innovation enables Mamba to achieve superior performance compared to other SSMs across various tasks.
Appendix B Broader Impacts
@@ C.1 Background on Graph Theory
We provide additional background on graph theory necessary for the definitions and theories of SETs and SENTs. The background is largely based on Diestel [17].We first give the formal definition of graph isomorphism:
@@
Graph isomorphism is an equivalence relation on graphs and as such it partitions the class of all graphs into equivalence classes. A set of graphs isomorphic to each other is called an isomorphism class of graphs. It is worth noting that our SENT isomorphism also partitions the class of all SENTs into equivalence classes in a similar fashion.We also provide the formal definition of induced subgraph:
@@ C.2 Remarks on Tokenized SENTs
[…]
@@ C.3 Remarks on Model Inference
[…]
@@ Appendix D Proofs
In this section, we provide proof for the theorems stated in the manuscript.
@@ E.1 Datasets
Both of these datasets are from Martinkus et al. [41]. The Planar dataset consists of 200 planar graphs with 64 nodes each, generated via Delaunay triangulation on points uniformly sampled in the unit square. The SBM dataset contains 200 graphs comprising 2 to 5 communities, with each community having between 20 and 40 nodes. An edge is placed between two nodes with probability 0.3 if they belong to the same community, and 0.05 otherwise. We follow the same splits as Martinkus et al. [41].
@@
The Proteins dataset includes graph representations (contact maps) of proteins from Dobson and Doig [18]. In these graphs, each node represents an amino acid, and an edge connects two nodes if their corresponding amino acids are within 6 angstroms of each other. We use the same data splits as Liao et al. [39]. The Point Clouds dataset, also from Liao et al. [39], consists of 41 point clouds of household objects [44]. As many of these graphs are disconnected, we retain only the largest connected component of each, following Bergmeister et al. [2], and again employ the splits used by Liao et al. [39].
@@
The QM9 dataset, from Wu et al. [63], comprises small molecules with up to nine heavy atoms (carbon, oxygen, nitrogen, and fluorine). In this work, we adopt the more challenging setting proposed by Vignac et al. [61], where hydrogen atoms are modeled explicitly, and we follow the same data splits as in that reference.
@@
The MOSES and GuacaMol datasets are obtained from the respective benchmark tools of Polykovskiy et al. [48] and Brown et al. [6]. Both consist of drug-like molecules, with those in GuacaMol typically being larger on average. For each dataset, we convert generated molecular graphs to SMILES using the code from Jo et al. [31], which permits partial charges. We employ the standard data splits provided by the corresponding benchmarks.
PubChem-10M.
PubChem-10M is a subset of about 10M molecules from PubChem curated by Chithrananda et al. [12].
@@
We generate the graphs using the generators from the NetworkX library [^2] https://networkx.org/documentation/stable/reference/generators.html [23], categories including “Classic”, “Lattice”, “Small”, “Random Graphs”, “Geometric”, “Trees”, “Community”, “Social Networks”. We ensure that this dataset does not contain any graphs in the downstream datasets. The summary of the code for generating these graphs is provided in Table 9. Notably, the largest graph has up to 5999 nodes.
@@ E.2 Evaluation Metrics
We follow Martinkus et al. [41] and Vignac et al. [61] in comparing our model’s performance with other graph generative approaches. Specifically, we measure the maximum mean discrepancy (MMD) between the generated and test graphs for degree distribution, clustering coefficient, orbit counts, and spectrum. As a reference, we also compute these metrics on the training set and report the mean ratio across all properties as a global indicator of statistical discrepancy between the generated samples and test samples. Note that for the Point Clouds dataset, which is defined by a $k$-nearest-neighbor structure, the degree MMD is always zero and is therefore excluded from the mean ratio. While we utilize these metrics to maintain consistency with previous research, we acknowledge their limitations, particularly regarding arbitrary kernel hyperparameter selection, as highlighted by O’Bray et al. [47], Thompson et al. [55]. In short, MMD measures the distributional similarity between generated and real graphs. Lower MMD scores indicate that the generated graphs’ statistics (e.g., degree, clustering coefficients) more closely match the training data.We additionally track uniqueness and novelty: uniqueness is the fraction of generated graphs that are not isomorphic to each other, and novelty is the fraction of generated graphs that are not isomorphic to any training graph.Below, we describe additional metrics specific to each dataset.
@@
Following Martinkus et al. [41], we report a validity score for synthetic datasets. For Planar graphs, it verifies whether the generated graphs remain planar; for SBM graphs, it measures how likely they are to be generated under the original SBM parameters. We integrate validity, novelty, and uniqueness into a single metric, VUN, which measures the fraction of generated graphs that are simultaneously valid, novel, and unique. In short, VUN assesses sample quality.
@@
For QM9, we report the validity, uniqueness, and novelty defined for general molecules, as described in the following paragraph.We also report atom stability and molecule stability as defined by Hoogeboom et al. [24] and Vignac et al. [61].
@@
Since MOSES [48] and GuacaMol [6] are benchmarking platforms, each comes with its own suite of metrics, which we use to evaluate our model. These include:
@@ E.4 Hyperparameters
Unlike prior studies that adjust model sizes across datasets, we maintain a consistent model architecture and size throughout all experiments, specifically using the small GPT configuration (768 hidden dimensions, 12 layers, 12 attention heads). Training hyperparameters are aligned with established practices from popular LLMs such as GPT-3 [7] and LLaMA [58]. We fix the context length to 2048 and use a batch size of 128 if possible, otherwise 64 for larger graphs. In particular, we employ the AdamW optimizer with a gradient clipping threshold of 1.0, a weight decay of 0.1, and a learning rate schedule with a linear warmup followed by cosine decay, peaking at 6e-4. The AdamW hyperparameters are set to $\beta=(0.9,0.95)$. Due to the small dataset sizes of previous benchmarks, we tune the only training hyperparameter dropout in $\{0,0.5\}$, and find the model achieves better validation loss with the value of 0.5 on the small synthetic datasets. Each model was trained for 200000, 400000, or 800000 iterations, depending on the dataset size.Inference hyperparameters, including $k$ (top-k sampling) and $\tau$ (temperature), are reported in Table 10 and analyzed in detail in Section F.5.